OpenStudy (a1234):

Proportion: (2+√2)/(2-√2) = x/(3+√10)

hero (hero):

\[(3 + \sqrt{10}) \dot\ \frac{2 + \sqrt{2}}{2 - \sqrt{2}}= x\]

OpenStudy (a1234):

Shouldn't we cross-multiply? With (3+√10) * (2+√2)?

OpenStudy (a1234):

That's the part that's confusing me.

hero (hero):

Yes, you should

OpenStudy (a1234):

So, how did you get what you put above?

hero (hero):

I Multiplied both sides by \(3 + \sqrt{10}\)

OpenStudy (a1234):

I'll try to show what I'm confused about...

OpenStudy (a1234):

This is what I thought it would be: \[x = ((3+\sqrt{10})(2 + \sqrt{2}))/(2 - \sqrt{2})\] So I don't get how we multiply both of (2+√2) / (2-√2) by (3+√10).

hero (hero):

What you have is correct. You only multiply it with the numerator.

OpenStudy (a1234):

Oh okay...so how do we proceed/

hero (hero):

Think of it this way \[\frac{a}{b} \times \frac{c}{d} = \frac{ab}{cd}\] So \[\frac{3 + \sqrt{10}}{1} \dot\ \frac{2 + \sqrt{2}}{2 - \sqrt{2}}= \frac{(3 + \sqrt{10})(2 + \sqrt{2})}{2 - \sqrt{2}}\]

OpenStudy (a1234):

Yes, I get that.

hero (hero):

To multiply \((3 + \sqrt{10})(2 + \sqrt{2})\) use the distributive property

OpenStudy (a1234):

So, would that be: (6 + 3√2 + 2√10 + √20)?

hero (hero):

Looks right

OpenStudy (a1234):

Alright, can you help me in dividing that by (2-√2)?

OpenStudy (a1234):

\[(6 + 3√2 + 2√10 + √20)\div (2-√2)\]

hero (hero):

Hang on

hero (hero):

Multiply the left side of the proportion by the conjugate first \[\frac{(2+√2)}{(2-√2)} \times \frac{2 + \sqrt{2}}{2 + \sqrt{2}} = \frac{x}{(3+√10)}\]

hero (hero):

Then simplify

OpenStudy (a1234):

Would that be 3 = x/(3x + √10)?

hero (hero):

You have to multiply it all out properly.

hero (hero):

Use FOIL to multiply if you have to

OpenStudy (a1234):

I'm getting 1 now, if I did it correctly.

hero (hero):

No, you did not do it correctly. Post your steps here so I can point out what you are doing wrong.

OpenStudy (a1234):

(2 + √2)(2 + √2) 4 + 2√2 + 2√2 + √4 4 + 4√2 + 2 6 + 4√2 (2 - √2)(2 + √2) 4 + 2√2 - 2√2 - √4 4 - 2 = 2 (6 + 4√2)/2

hero (hero):

That's better, but you're not done simplifying

OpenStudy (a1234):

Would it be 3 + 2√2?

hero (hero):

Yes, now multiply both sides by \(3 + \sqrt{10}\) to isolate \(x\)

OpenStudy (a1234):

(3 + √10)(3 + 2√2) 9 + 6√2 + 3√10 + 2√20

hero (hero):

Yes, but now you have to simplify \(2\sqrt{20}\)

OpenStudy (a1234):

9 + 6√2 + 3√10 + 2√20 2√ 4 * 5 4√5 9 + 6√2 + 3√10 + 4√5

hero (hero):

Congrats

OpenStudy (a1234):

Thank you so much!

hero (hero):

yw